OpenStudy (anonymous):

Simplify the expression (2√27) - (3√32).

OpenStudy (anonymous):

\[6\sqrt{3}-12\sqrt{2}\]

OpenStudy (lukecrayonz):

27 can be broken down into 9*3, and nine can be broken down into 3*3. It's a square root, so you will look for PAIRS of common terms, and you have three 3's. You can remove two, and put them on the outside. The remaining will stay inside. So 2*3sqrt(3), also 6sqrt(3).

OpenStudy (lukecrayonz):

32, 16*2, and 16=4*4. Remove the fours, and bring them outside, for 3*4sqrt(2), also 12sqrt(2)

OpenStudy (anonymous):

i'm sorry i think that should be multiplied not subtracted

OpenStudy (anonymous):

hard to read the problem in the book

OpenStudy (lukecrayonz):

Well just take what I made, (6sqrt(3))(12sqrt(2)), then multiply. 6*12=72, 3*2=6, so your ending answer is 72sqrt(6)

OpenStudy (anonymous):

thanks